# JLQCD collaboration

### v2

@@ I Introduction
In the massless limit of the up and down quarks, it is natural to assume that the standard $SU(2)_{L}\times SU(2)_{R}$ symmetry is recovered at some critical temperature in the range 150–200 MeV and no other symmetry exists. However, the axial $U(1)_{A}$, which is broken by anomaly, may effectively appear if nontrivial topological excitations of gluons are strongly suppressed above $T_{c}$. When such topological effects are described by instantons Gross:1980br ; Diakonov:1984vw ; Morris:1984zi ; Schafer:1996wv ; Ringwald:1999ze , which is valid for pure Yang-Mills, QCD with heavy quarks, and QCD with a large number of color degrees of freedom, strong suppression of the quantum anomaly is not expected and the axial $U(1)_{A}$ is likely to remain broken at $T\sim T_{c}$. However, with the light dynamical quarks, the typical size of the topological excitation is described by the pion physics whose correlation length becomes larger than $1/T$,and this description of the topological excitations by instantons is no longer valid. The fate of the axial $U(1)_{A}$ anomaly at high $T$ has been discussed in both theoretical works Cohen:1996ng ; Cohen:1997hz ; Aoki:2012yj ; Pelissetto:2013hqa ; Kanazawa:2014cua ; Sato:2014axa ; Nakayama:2016jhq as well as numerical works Cossu:2013uua ; Buchoff:2013nra ; Dick:2015twa ; Brandt:2016daq ; Ishikawa:2017nwl ; Bazavov:2019www .
@@
Another interesting possibility is that at high $T$, QCD may in fact have a larger symmetry. This symmetry is related to the anti-periodic boundary condition of quarks imposed in the temporal direction. At very high temperature, the mesons would have large masses proportional to the temperature $T$, due to the Matsubara mass. As a consequence, a nontrivial symmetry may appear, analogous to the heavy quark symmetry Neubert:1993mb ; Brambilla:2004jw at $T=0$. In fact, in the perturbative evaluation in Laine:2003bd the masses of all the mesons of different spins converge to $2\pi T$. Note that this is not a symmetry of the original QCD Lagrangian but is an “emergent” symmetry in the large $T$ limit. Although this is only an approximate symmetry within the perturbative framework in the large $T$ limit, nonperturbative gluonic effects may enhance this symmetry even at $T\sim 2T_{c}$, this symmetry is called the chiral-spin symmetry Glozman:2014mka ; Glozman:2016swy ; Glozman:2017dfd ; Lang:2018vuu ; Catillo:2021rrq ; Glozman:2022lda .
@@
It is, therefore, a crucial task to investigate the symmetry at high temperatures through a nonperturbative approach to QCD, which must include a consistency check with the perturbative prediction of the screening masses converging to $2\pi T$. Lattice QCD simulations that have previously studied this issue, however, used fermion formulations that explicitly broke the $SU(2)_{L}\times SU(2)_{R}$ symmetry. This has made the analysis ambiguous in identifying the symmetry breaking/restoration patterns apart from lattice artifacts. In particular, we reported in Cossu:2015kfa ; Tomiya:2016jwr ; Aoki:2020noz that the lattice artifacts in the probes of axial $U(1)_{A}$ anomaly are enhanced at high temperatures and it is difficult to disentangle the physical effect from the discretization errors.
@@
In this work, we employ the domain-wall fermion formalism Kaplan:1992bt ; Shamir:1993zy ; Furman:1994ky with an improvement proposed in Refs.Brower:2005qw ; Brower:2012vk . With this Möbius domain-wall fermion, which is a good approximation of the overlap fermion Neuberger:1997fp , we achieve a theoretically clean control of the symmetry of QCD. With a fixed lattice spacing $a\sim 0.075$ fm, the residual mass, which is an indicator of the explicit symmetry violation is reduced to 0.14(6) MeV. In addition, we set the quark masses so that the lightest bare quark mass $ma=0.001$ is less than the physical mass value of the up and down quarks. The simulated temperatures span a range $T\in[147,330]$ MeV, which corresponds to $[0.9,2]T_{c}$, where the $T_{c}$ is estimated from the peak of the chiral susceptibility Aoki:2024uvl . The main runs are carried out with a fixed lattice size $L=32$ except for the lowest two temperatures, where the pions behave as the pseudo-Nambu-Goldstone bosons; in this case we simulate two additional volumes, $L=40$ and $48$, to control the finite volume systematics.
@@
On each ensemble of O(1000-10000) trajectories, we measure the two-point mesonic correlation functions in 6 different iso-triplet channels every 50 trajectories. With the screening masses obtained from fitting the correlators, we examine the symmetry at each simulated temperature and quark mass. In our main analysis of the mesonic correlators, the standard $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry is examined by the difference between the vector $V$ and axial vector $A$ correlators. The tensor $T_{t}$ and $X_{t}$ channels (definitions will be given below), rather than scalar $S$(noisy) and pseudoscalar $PS$ channels, are useful to estimate how much the axial $U(1)_{A}$ anomaly effect remains.
Because of the Matsubara mass, the system may exhibit the chiral-spin $SU(2)_{CS}$ symmetry mentioned above. In our previous work at higher temperatures Rohrhofer:2017grg ; Rohrhofer:2019qwq ; Rohrhofer:2019qal we observed the emergence of this $SU(2)_{CS}$ symmetry from the correlator ratios among $V$, $A$, $T_{t}$ and $X_{t}$. A similar observation simulating quarks with a good chiral symmetry was reported in Refs. Chiu:2023hnm ; Chiu:2024jyz ; Chiu:2024bqx . In this work, we revisit this enlarged symmetry group, computing the screening masses and simulating lower temperatures down to $T\sim 147$ MeV, which is $\sim 0.9T_{c}$.
@@ II Mesonic two-point functions and symmetries at high temperatures
We do not distinguish $k=1$ and $k=2$, since they are identical under the rotational symmetry along the $z$ axis. The generators $\Sigma^{i}$ form an $SU(2)$ algebra, which is identical to the chiral-spin $SU(2)_{CS}$ symmetry suggested in the literature Glozman:2014mka ; Glozman:2016swy ; Glozman:2017dfd ; Lang:2018vuu ; Catillo:2021rrq ; Glozman:2022lda .
Note that any correction from gauge field potential only appears in the next-to-leading $1/T$ contribution. Thus, the chiral spin $SU(2)_{CS}$ symmetry emerges at sufficiently high temperatures. With this $SU(2)_{CS}$ symmetry, the mesonic operators $(A_{1},T_{t},X_{t})$ form a nontrivial triplet. These operators have different spins in the original four dimensions, but they are all in the spin-one representation of $SO(3)$ symmetry along the $z$ axis. One can consider another set $(V_{2},PS,S)$ but the operators transform differently under the $SO(3)$ rotation and therefore, the mass degeneracy would not be that good111 We thank L. Glozman for pointing out the difference of the two triplets..
@@
with a numerical constant $c$ coming from the one-loop correction where $g$ is the gauge coupling. It is remarkable that the result does not depend on $\Gamma$. This result is analogous to the heavy quark symmetry Neubert:1993mb ; Brambilla:2004jw which is an approximate symmetry of the bottom or charm quarks at $T=0$. As the spin and orbital angular momentum dependent interactions are proportional to the inverse of the heavy quark mass, the spectrum of the quarkonia and heavy-light mesons are insensitive to spin. The same may explain here that with a large Matsubara mass proportional to $T$, the operators $(V_{1},PS,S)$ with different spin form an $SU(2)_{CS}$ triplet222 The $SO(3)$ rotational symmetry is broken also among $(A_{1},T_{t},X_{t})$ at very high temperatures, since the temporal direction shrinks as $1/T\to 0$. However, its effect is sub-leading. . In Ref. Chiu:2023hnm , it was suggested that other combinations of tensors $\gamma_{1}\gamma_{3}(T_{k})$ and $\gamma_{1}\gamma_{3}\gamma_{5}(X_{k})$ form a quartet with $A_{4}$ and $V_{4}$. Inclusion of this quartet into our analysis is of interest to future work; however, in this paper we will focus on the triplet.
It is an important numerical subject to quantify the emergent $SU(2)_{CS}$ symmetry at intermediate temperatures $T\sim T_{c}$ and see if it has any impact on the chiral phase transition. It is also interesting to investigate if an additional transition exists at $T>T_{c}$, triggered by this emergent symmetry. These three symmetries, accessible through the differences between mesonic screening masses from the pairs of the triplet channels333 In our previous study Rohrhofer:2019qal , we also studied the ratio of differences of correlators $|C_{A}(z)-C_{T}(z)|/|C_{A}(z)-C_{S}(z)|$. However, we find that the denominator approaches to zero at high temperatures (presumably due to the large Matsubara mass and the analogy to “heavy quark symmetry” we discussed in Sec. II.) and the ratio is not a good indicator of the $SU(2)_{CS}$ symmetry. In this work, we focus on the screening mass differences and present the values in the physical unit (MeV) so that one can easily compare with the simulated temperature for each set of data. , constitute our motivation to simulate QCD at high temperatures in the range $[0.9,2]T_{c}$, for a range of quark masses covering the physical point as a method to understand the symmetry structure of QCD at high temperatures.
@@ III Lattice setups
The lightest bare sea quark mass $m=0.001$ corresponds to 2.6 MeV, which is estimated to be 71% of the physical quark mass $m_{\text{phys}}=0.0014(2)$.
For the link variables in the Dirac operator, we perform the stout smearing Morningstar:2003gk three times with the smoothing parameter $\rho=0.1$. The size of the extra dimension for the Möbius domain-wall fermion action is set to $L_{s}=16$. With this treatment, the residual mass or the chiral symmetry breaking of the lattice Dirac operator is suppressed to 0.14(6) MeV.
@@ IV Numerical results
In Fig. 7, we plot the mass difference between $A$ and $V$ channels $\Delta m_{V-A}$ , which probes the $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry breaking, as a function of temperature $T$. Here the data from the largest spatial volume for each simulated temperature is plotted. Around the critical temperature shown by the vertical grey band, the mass difference $\Delta m_{V-A}$ quickly vanishes. This is most clearly seen at our lightest mass $m=0.001$, we find $\Delta m_{V-A}/m_{A}\sim 0.1$% at $T=190$ MeV and 2.8% at $T=165$ MeV666 In recent study Gavai:2024mcj with Möbius domain-wall quarks, violation of the axial $U(1)_{A}$ symmetry even at $T=186$ MeV was reported. We notice, however, that their lattice spacings $a>0.13$ fm are coarser than the one $a\sim 0.1$ fm we simulated in our previous work Cossu:2015kfa where we find a larger violation of the Ginsparg-Wilson in the low-lying eigenmodes of the Dirac operator than the residual mass, which leads to overestimation of the axial $U(1)_{A}$ anomaly. . Both values are consistent with zero to within a standard deviation.
@@
A similar behavior but with larger errorbars is found in Figs. 7 and 9 where the mass differences between $X_{t}$ and $T_{t}$, $\Delta m_{X-T}$, and that of $PS$ and $S$, $\Delta m_{PS-S}$, are plotted. These are the probes for the axial $U(1)_{A}$ symmetry. We find that $\Delta m_{X-T}/m_{X}$ is 2.5% at $T=190$ MeV and 12.7% at $T=165$ MeV. The latter is consistent with zero within a larger errorbar. This indicates that the chiral anomaly effect is quite suppressed above $T_{c}$.
@@
Fig. 9 is a plot of the chiral-spin $SU(2)_{CS}$ symmetry, which is an approximate symmetry expected up to $O(1/T)$ corrections, as explained in Sec. II. This symmetry is examined by the mass difference $\Delta m_{X-A}$ between $X_{t}$ and $A$. The $T$ dependence for the screening mass, in addition to the reduction in noise, is analogous to $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$; however, we find a qualitative difference of $SU(2)_{CS}$ in Fig. 10. Here we plot the quark mass dependence of $\Delta m_{V-A}$ (squares), $\Delta m_{PS-S}$ (circles), $\Delta m_{X-T}$ (upward triangles), and $\Delta m_{X-A}$ (downward triangles) at $T=330$ MeV. The former three probing $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries, are consistent with zero, and their uncertainty is $\sim 1$MeV, which is less than 1% of $T$. In contrast to these other mass differences, the data $\Delta m_{X-A}$ is clearly nonzero even at the lightest mass: $\sim-40$ MeV showing that the associated chiral-spin symmetry $SU(2)_{CS}$ is only approximate: $|\Delta m_{X-A}|/T\sim 0.12$. We note that at much higher temperatures than $T=1$ GeV, a better degeneracy even between $PS$ and $V$ channels was reported in Ref. DallaBrida:2021ddx .
@@
Figure 1: Effective screening mass plots at $T=147$ MeV. The data for $PS,S,A,T_{t}$ channels at the lightest quark mass $m=0.001$ are presented.
@@
Figure 9: Temperature $T$ vs. the screening mass difference between $A_{x}$ and $X_{t}$ which probes the $SU(2)_{CS}$ symmetry. The vertical grey band indicates our estimate for the critical temperature $T_{c}=165(3)$ MeV from the chiral susceptibility.
Figure 10: Quark mass dependence of $\Delta m_{V-A}$ (squares), $\Delta m_{PS-S}$ (circles), $\Delta m_{X-T}$ (upward triangles), and $\Delta m_{X-A}$ at $T=330$ MeV. Compared to the former three probing $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries, which are consistent with zero with errorrbars $\sim 1$ MeV, the chiral-spin probe $\Delta m_{X-A}$ (downward triangles) is clearly nonzero: $\sim-40$ MeV and the associated chiral-spin symmetry $SU(2)_{CS}$ is only approximate: $|\Delta m_{X-A}|/T\sim 0.12$.

### v4

@@ I Introduction
In the massless limit of the up and down quarks, it is natural to assume that the standard $SU(2)_{L}\times SU(2)_{R}$ symmetry is recovered at some critical temperature in the range 150–200 MeV and no other symmetry exists. However, the axial $U(1)_{A}$, which is broken by anomaly, may effectively appear if nontrivial topological excitations of gluons are strongly suppressed above $T_{c}$. When such topological effects are described by instantons Gross:1980br ; Diakonov:1984vw ; Morris:1984zi ; Schafer:1996wv ; Ringwald:1999ze , which is valid for pure Yang-Mills, QCD with heavy quarks, and QCD with a large number of color degrees of freedom, strong suppression of the quantum anomaly is not expected and the axial $U(1)_{A}$ is likely to remain broken at $T\sim T_{c}$. However, with the light dynamical quarks, the typical size of the topological excitation is described by the pion physics whose correlation length becomes larger than $1/T$, and this description of the topological excitations by instantons is no longer valid. The fate of the axial $U(1)_{A}$ anomaly at high $T$ has been discussed in both theoretical works Cohen:1996ng ; Cohen:1997hz ; Aoki:2012yj ; Pelissetto:2013hqa ; Kanazawa:2014cua ; Sato:2014axa ; Nakayama:2016jhq as well as numerical works Cossu:2013uua ; Buchoff:2013nra ; Dick:2015twa ; Brandt:2016daq ; Ishikawa:2017nwl ; Bazavov:2019www ; Kaczmarek:2023bxb .
@@
Another interesting possibility is that at high $T$, QCD may in fact have a larger symmetry. This symmetry is related to the anti-periodic boundary condition of quarks imposed in the temporal direction. At very high temperature, the mesons would have large masses proportional to the temperature $T$, due to the Matsubara mass. In fact, in the perturbative evaluation in Laine:2003bd the masses of all the mesons of different spins converge to $2\pi T$ at lowest order. From this point of view, we could consider the light quarks which form the meson to become sufficiently “heavy” at high temperatures leading to a spin independent symmetry. Such a symmetry is analogous to the heavy quark symmetry Neubert:1993mb ; Brambilla:2004jw in $T=0$ QCD. Note that this is not a symmetry of the original QCD Lagrangian but is an “emergent” symmetry in the large $T$ limit. Although this is only an approximate symmetry within the perturbative framework in the large $T$ limit, nonperturbative gluonic effects may enhance this symmetry even at $T\sim 2T_{c}$, this symmetry is called the chiral-spin symmetry Glozman:2014mka ; Glozman:2016swy ; Glozman:2017dfd ; Lang:2018vuu ; Catillo:2021rrq ; Glozman:2022lda .
@@
Our goal in this work is to simulate QCD with chiral fermions and to investigate these symmetries at high temperatures, including how the meson masses approach the perturbative prediction $2\pi T$111 In DallaBrida:2021ddx , they simulated QCD at very high temperatures $T\geq 1$ GeV and found that the meson masses approach $2\pi T$ from above, which is consistent with Laine:2003bd although the channel-dependence is still visible. . Lattice QCD simulations that have previously studied this issue, however, used fermion formulations that explicitly broke the $SU(2)_{L}\times SU(2)_{R}$ symmetry. For the Wilson fermion, the symmetry is broken down to a vectorlike $SU(2)_{V}$ and for the staggered fermion, only the $U(1)_{V^{\prime}}\times U(1)_{A^{\prime}}$ subgroup (which are different from the standard quark number and anomalous axial $U(1)_{A}$ symmetries) remains. In particular, we reported in Cossu:2015kfa ; Tomiya:2016jwr ; Aoki:2020noz that the lattice artifacts in the probes of axial $U(1)_{A}$ anomaly are enhanced at high temperatures and it is difficult to disentangle the physical effect from the discretization errors.
@@
In this work, we employ the domain-wall fermion formalism Kaplan:1992bt ; Shamir:1993zy ; Furman:1994ky with an improvement proposed in Refs.Brower:2005qw ; Brower:2012vk . With this Möbius domain-wall fermion, which is a good approximation of the overlap fermion Neuberger:1997fp , we achieve a theoretically clean control of the symmetry of QCD. With a fixed lattice spacing $a\sim 0.075$ fm, the residual mass Tomiya:2016jwr , which is an indicator of the explicit symmetry violation is reduced to 0.14(6) MeV. In addition, we set the quark masses so that the lightest bare quark mass $ma=0.001$ is less than the physical mass value of the up and down quarks. The simulated temperatures span a range $T\in[147,330]$ MeV, which corresponds to $[0.9,2]T_{c}$, where the $T_{c}$ is estimated from the peak of the chiral susceptibility Aoki:2024uvl . The main runs are carried out with a fixed lattice size $L=32$ except for the lowest two temperatures, where the pions behave as the pseudo-Nambu-Goldstone bosons; in this case we simulate two additional volumes, $L=40$ and $48$, to control the finite volume systematics. While this study is carried out for a fixed lattice spacing, our lattice spacing $a\sim 0.075$ fm is at the same level of the finest lattice in the literature. In addition to this, our use of the Möbius domain wall fermion introduces an automatic $\mathcal{O}(a)$ improvement over dicretization effects; while these cannot be estimated in this study, future work will be to make an extrapolation to the continuum limit.
@@
On each ensemble of O(1000-10000) trajectories, we measure the two-point mesonic correlation functions in 6 different iso-triplet channels every 50 trajectories. With the screening masses obtained from fitting the correlators, we examine the symmetry at each simulated temperature and quark mass. In our main analysis of the mesonic correlators, the standard $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry is examined by the difference between the vector $V$ and axial vector $A$ correlators. The tensor $T_{t}$ and $X_{t}$ channels (definitions will be given below), rather than scalar $S$(noisy) and pseudoscalar $PS$ channels, are useful to estimate how much the axial $U(1)_{A}$ anomaly effect remains. In lattice QCD, the isospin triplet $S$ channel has been difficult to precisely measure. As a possible reason, it was pointed out in Mathur:2006bs that the isospin triplet $S$ channel may couple to many composite meson states. Therefore, study of the $U(1)_{A}$ as well as any other underlying quark symmetries using the $S$ channel should be considered alongside an additional probe. Because of the Matsubara mass, the system may exhibit the chiral-spin $SU(2)_{CS}$ symmetry mentioned above. In our previous work at higher temperatures Rohrhofer:2017grg ; Rohrhofer:2019qwq ; Rohrhofer:2019qal we observed the emergence of this $SU(2)_{CS}$ symmetry from the correlator ratios among $V$, $A$, $T_{t}$ and $X_{t}$. A similar observation simulating quarks with a good chiral symmetry was reported in Refs. Chiu:2023hnm ; Chiu:2024jyz ; Chiu:2024bqx . In this work, we revisit this enlarged symmetry group, computing the screening masses and simulating lower temperatures down to $T\sim 147$ MeV, which is $\sim 0.9T_{c}$.
@@ II Mesonic two-point functions and symmetries at high temperatures
We do not distinguish $k=1$ and $k=2$, since they are identical under the rotational symmetry along the $z$ axis. The generators $\Sigma^{i}$ form an $SU(2)$ algebra, which is identical to the chiral-spin $SU(2)_{CS}$ symmetry suggested in the literature Glozman:2014mka ; Glozman:2016swy ; Glozman:2017dfd ; Lang:2018vuu ; Catillo:2021rrq ; Glozman:2022lda 222 In the hadron correlators, the relative momentum part in $p_{1}$ and $p_{2}$ must be integrated out but that ends up with a typical scale $1/z$ and the correction is $1/zT$ as clearly shown in Rohrhofer:2019qwq . .
Note that any correction from gauge field potential only appears in the next-to-leading $1/T$ contribution. Thus, the chiral spin $SU(2)_{CS}$ symmetry emerges at sufficiently high temperatures. With this $SU(2)_{CS}$ symmetry, the mesonic operators $(A_{1},T_{t},X_{t})$ form a nontrivial triplet. These operators have different spins in the original four dimensions, but they are all in the spin-one representation of $SO(3)$ symmetry along the $z$ axis. One can consider another set $(V_{2},PS,S)$ but the operators transform differently under the $SO(3)$ rotation and therefore, the mass degeneracy would not be that good.
@@
with a numerical constant $c$ coming from the one-loop correction where $g$ is the gauge coupling. It is remarkable that the result does not depend on $\Gamma$. This result is analogous to the heavy quark symmetry Neubert:1993mb ; Brambilla:2004jw which is an approximate symmetry of the bottom or charm quarks at $T=0$. As the spin and orbital angular momentum dependent interactions are proportional to the inverse of the heavy quark mass, the spectrum of the quarkonia and heavy-light mesons are insensitive to spin. The same may explain here that with a large Matsubara mass proportional to $T$, the operators $(V_{1},PS,S)$ with different spin form an $SU(2)_{CS}$ triplet333 The $SO(3)$ rotational symmetry is broken also among $(A_{1},T_{t},X_{t})$ at very high temperatures, since the temporal direction shrinks as $1/T\to 0$. However, its effect is sub-leading. . In Ref. Chiu:2023hnm , it was suggested that other combinations of tensors $\gamma_{1}\gamma_{3}(T_{k})$ and $\gamma_{1}\gamma_{3}\gamma_{5}(X_{k})$ form a quartet with $A_{4}$ and $V_{4}$. Inclusion of this quartet into our analysis is of interest to future work; however, in this paper we will focus on the triplet.
It is an important numerical subject to quantify the emergent $SU(2)_{CS}$ symmetry at intermediate temperatures $T\sim T_{c}$ and see if it has any impact on the chiral phase transition. It is also interesting to investigate if an additional transition exists at $T>T_{c}$, triggered by this emergent symmetry. These three symmetries, accessible through the differences between mesonic screening masses from the pairs of the triplet channels444 In our previous study Rohrhofer:2019qal , we also studied the ratio of differences of correlators $|C_{A}(z)-C_{T}(z)|/|C_{A}(z)-C_{S}(z)|$. However, we find that the denominator approaches to zero at high temperatures (presumably due to the large Matsubara mass and the analogy to “heavy quark symmetry” we discussed in Sec. II.) and the ratio is not a good indicator of the $SU(2)_{CS}$ symmetry. In this work, we focus on the screening mass differences and present the values in the physical unit (MeV) so that one can easily compare with the simulated temperature for each set of data. , constitute our motivation to simulate QCD at high temperatures in the range $[0.9,2]T_{c}$, for a range of quark masses covering the physical point as a method to understand the symmetry structure of QCD at high temperatures.
@@ III Lattice setups
We take 4–6 different quark masses depending on the temperatures. The lightest bare sea quark mass $m=0.001$ corresponds to 2.6 MeV, which is estimated to be 71% of the physical quark mass $m_{\text{phys}}=0.0014(2)$. In this study, we always take the valence and sea quark masses equal.
For the link variables in the Dirac operator, we perform the stout smearing Morningstar:2003gk three times with the smoothing parameter $\rho=0.1$. The size of the extra dimension for the Möbius domain-wall fermion action is set to $L_{s}=16$. With this treatment, the residual mass or the chiral symmetry breaking of the lattice Dirac operator is suppressed to $0.14(6)$ MeV Tomiya:2016jwr .
@@ IV Numerical results
To obtain the screening masses from the correlation functions the effective mass is extracted from data by applying Newton’s method to the cosh ansatz form of the spatial correlator. We then perform an uncorrelated least square fit by finding a suitable range for which the effective mass does not exceed $2\sigma$ from the error band. Regions where the error became 100% are omitted from the fits where possible. When the correlators are stable, the channels connected by symmetry transformations e.g. $V-A$, $PS-S$ and $X-T$ are fitted to the same range. We also confirm that the screening mass differences do not depend on the fitting range very much and the systematics are well controlled compared to the statistical uncertainties.
We find that the $S$ channel correlators are quite unstable, reflecting exceptionally large contribution from a few configurations. We decide to discard the $S$ data when 1)the correlator goes negative or 2)there is no stable plateau in the fitting range of the symmetry partner $PS$.
@@
In Fig. 7, we plot the mass difference between $A$ and $V$ channels $\Delta m_{V-A}$ , which probes the $SU(2)_{L}\times SU(2)_{R}$ chiral symmetry breaking, as a function of temperature $T$. Here the data from the largest spatial volume for each simulated temperature is plotted. Around the critical temperature shown by the vertical grey band, the mass difference $\Delta m_{V-A}$ quickly vanishes. This is most clearly seen at our lightest mass $m=0.001$, we find $\Delta m_{V-A}/m_{A}\sim 0.1$% at $T=190$ MeV and 2.8% at $T=165$ MeV. Further evidence of this rapid drop off is also seen in the literature with a rapid suppression of several order parameters associated with chiral symmetry shown in Krasniqi:2024kwm . For the higher quark masses, $\Delta m_{V-A}$ shows a milder $T$-dependence, which suggests that the crossover temperature becomes higher. Both values are consistent with zero to within a standard deviation.
@@
A similar behavior but with larger errorbars is found in Figs. 7 and 9 where the mass differences between $X_{t}$ and $T_{t}$, $\Delta m_{X-T}$, and that of $PS$ and $S$, $\Delta m_{PS-S}$, are plotted. These are the probes for the axial $U(1)_{A}$ symmetry. We find that $\Delta m_{X-T}/m_{X}$ is 2.5% at $T=190$ MeV and 12.7% at $T=165$ MeV777 In recent study Gavai:2024mcj with Möbius domain-wall quarks, violation of the axial $U(1)_{A}$ symmetry even at $T=186$ MeV was reported. We notice, however, that their lattice spacings $a>0.13$ fm are coarser than the one $a\sim 0.1$ fm we simulated in our previous work Cossu:2015kfa where we find a larger violation of the Ginsparg-Wilson in the low-lying eigenmodes of the Dirac operator than the residual mass, which leads to overestimation of the axial $U(1)_{A}$ anomaly. . The latter is consistent with zero within a larger errorbar. This indicates that the chiral anomaly effect is quite suppressed above $T_{c}$.
@@
Fig. 9 is a plot of the chiral-spin $SU(2)_{CS}$ symmetry, which is an approximate symmetry expected up to $O(1/T)$ corrections, as explained in Sec. II. This symmetry is examined by the mass difference $\Delta m_{X-A}$ between $X_{t}$ and $A$. The $T$ dependence for the screening mass, in addition to the reduction in noise, is analogous to $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$; however, we find a qualitative difference of $SU(2)_{CS}$ in Fig. 10. Here we plot the quark mass dependence of $\Delta m_{V-A}$ (circles), $\Delta m_{PS-S}$ (squares), $\Delta m_{X-T}$ (upward triangles), and $\Delta m_{A-X}$ (downward triangles) at $T=330$ MeV. The former three probing $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries, are consistent with zero, and their uncertainty is $\sim 1$MeV, which is less than 1% of $T$. In contrast to these other mass differences, the data $\Delta m_{X-A}$ is clearly nonzero even at the lightest mass: $\sim-40$ MeV showing that the associated chiral-spin symmetry $SU(2)_{CS}$ is only approximate: $|\Delta m_{X-A}|/T\sim 0.12$. We note that at much higher temperatures than $T=1$ GeV, it was reported in Ref. DallaBrida:2021ddx that the mass difference between the $PS$ and $V$ channels goes down to $\sim 1$% of $2\pi T$ although, the difference is still statistically significant and deviates from the next-to-leading order perturbative QCD prediction, where the one-loop correction of each meson mass to $2\pi T$ is positive and channel-independent (for related results and discussion see Laudicina:2022ghk ; Giusti:2024ohu ).
@@
Figure 1: Effective screening mass plots at $T=147$ MeV. The data for $PS,S,A,X_{t}$ channels at the lightest quark mass $m=0.001$ are presented.
@@
Figure 9: Temperature $T$ vs. the screening mass difference between $A$ and $X_{t}$ which probes the $SU(2)_{CS}$ symmetry. The vertical grey band indicates our estimate for the critical temperature $T_{c}=165(3)$ MeV from the chiral susceptibility.
Figure 10: Quark mass dependence of $\Delta m_{V-A}$ (circles), $\Delta m_{PS-S}$ (squares), $\Delta m_{X-T}$ (upward triangles), and $\Delta m_{A-X}$(downward triangles) at $T=330$ MeV. Compared to the former three probing $SU(2)_{L}\times SU(2)_{R}$ and $U(1)_{A}$ symmetries, which are consistent with zero with errorrbars $\sim 1$ MeV, the chiral-spin probe $\Delta m_{A-X}$ is clearly nonzero: $\sim-40$ MeV and the associated chiral-spin symmetry $SU(2)_{CS}$ is only approximate: $|\Delta m_{X-A}|/T\sim 0.12$.